Assume that when adults with smartphones are randomly​ selected, 44% use them in meetings or classes. If 7 adult smartphone users are randomly​ selected, find the probability that exactly 5 of them use their smartphones in meetings or classes.

The probability is

(round to four decimals places as needed.)
Answer #1

Solution

Given that ,

p = 0.44

1 - p = 1 - 0.44 = 0.56

n = 7

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 5) = (7 C 5) * (0.44)5 * (0.56)2

= 0.108607

Probability = 0.1086